# Nonlinear Multi-objective Reinforcement Learning with Provable Guarantees

### v2

@@ Abstract
We describe RA-E3 (Reward-Aware Explicit Explore or Exploit), an algorithm with provable guarantees for solving a single or multi-objective Markov Decision Process (MDP) where we want to maximize the expected value of a nonlinear function over accumulated rewards. This allows us to model fairness-aware welfare optimization for multi-objective reinforcement learning as well as risk-aware reinforcement learning with nonlinear Von Neumann-Morgenstern utility functions in the single objective setting. RA-E3 extends the classic E3 algorithm that solves MDPs with scalar rewards and linear preferences. We first state a distinct reward-aware version of value iteration that calculates a non-stationary policy that is approximately optimal for a given model of the environment. This sub-procedure is based on an extended form of Bellman optimality for nonlinear optimization that explicitly considers time and current accumulated reward. We then describe how to use this optimization procedure in a larger algorithm that must simultaneously learn a model of the environment. The algorithm learns an approximately optimal policy in time that depends polynomially on the MDP size, desired approximation, and smoothness of the nonlinear function, and exponentially on the number of objectives.
@@ 1 Introduction
Markov Decision Processes (MDPs) model goal-driven interaction with a stochastic environment. The standard problem formulation specifies a scalar-valued reward per time-step and requires an agent to learn a policy (a mapping of states to actions) that maximizes expected (typically time-discounted) total reward. Equivalently, this formulation asks the agent to maximize the expected value of a linear function of total reward. This problem can be solved with provable approximation guarantees in polynomial time with respect to the size of the MDP (Kearns and Singh 2002; Brafman and Tennenholtz 2003; Auer, Jaksch, and Ortner 2008; Azar, Osband, and Munos 2017; Agarwal et al. 2021).
We study the generalization in which we want to optimize a nonlinear function of total reward, where the reward may be vector-valued to represent a multi-objective MDP (MOMDP). We call such a function the welfare function $W$, and we seek to maximize expected welfare $\mathbb{E}\left[W(\mathbf{r})\right]$ where $\mathbf{r}$ is a total reward vector $\mathbf{r}\in\mathbb{R}^{d}$ for a MOMDP with $d$ objectives and the expectation is with respect to randomness in the environment and (possibly) the learned policy. We note that this function is also canonically called the scalarization function within the multiobjective optimization literature (Hayes et al. 2023). We point out at the outset that this nonlinearity is entirely in the objective, as opposed to the many recent works on using nonlinear function approximation as a technique in the optimization of a linear objective over large state spaces.
@@
Nonlinear welfare functions allow us to express a richer space of preferences for the learning agent. For example, in computational social choice, one may express a desire for fairness or balance across objectives by selecting a welfare functions that exhibits diminishing marginal returns such as the Nash Social Welfare $\left(W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}\right)$ (Caragiannis et al. 2019). Even with a single objective, classical decision theory formulates differing risk-attitudes as the maximization of a potentially nonlinear function, also known as a Von Neumann-Morgenstern utility function (Von Neumann and Morgenstern 1947). For example, maximizing the expected value of a concave function of total reward like the logarithm models risk aversion, whereby one prefers a policy that always obtains reward $x$ over a policy that with 50% probability obtains reward $2x$ and otherwise obtains 00. The classic approach of optimizing total scalar reward expresses risk-neutrality, under which the learning agent is indifferent between these two policies.
Contributions and Outline
In this work, we describe the RA-E3 algorithm (Reward-Aware Explicit Explore or Exploit) for solving the problem of maximizing the expected value of a nonlinear function in an MOMDP. To the best of our knowledge, it is the first algorithm to provide provable worst-case approximation guarantees. In the following we summarize our contributions and outline the rest of the paper.
  • In Section 4 we expand the classic model of reinforcement learning to optimize the expected value of a nonlinear welfare function of (possibly) multiple dimensions of reward, motivated by principled multi-objective optimization and modeling of risk attitudes. Our major assumption on the welfare function is that it is smooth. We describe an extended form of Bellman optimality that describes the structure of optimal policies for this problem.
  • In Section 5 we describe Reward-Aware Value Iteration (RA-Value Iteration), a model-based algorithm for computing a non-stationary policy with approximately optimal expected welfare. The algorithm has runtime that is polynomially bounded in the size of the model, horizon time, smoothness of the welfare function, and approximation, and exponential in the number of objectives.
  • In Section 6 we describe how to remove the assumption of a model by incorporating RA-Value Iteration into the explore or exploit framework of the E3 algorithm  (Kearns and Singh 2002). We call the resulting overall algorithm Reward Aware Explicit Explore or Exploit, or RA-E3. In a number of steps bounded polynomially in all relevant parameters except the number of objectives, with high probability, the algorithm will halt in a state $s$ and output a policy $\pi$ that is approximately optimal with respect to the expected welfare beginning at $s$ with time and current reward reset to 0.
@@ 2 Related Work
Most algorithms for reinforcement learning focus on the case of a single scalar-valued objective where the goal is to maximize total expected reward (Sutton and Barto 2018). Multi-objective reinforcement learning (Liu, Xu, and Hu 2014) attempts to optimize multiple objectives at once. So-called single-policy methods use a scalarization function to reduce the problem to scalar optimization for a single policy, and we follow this line of research.
The simplest form is linear scalarization, applying a weighted sum on the Q vector (Moffaert, Drugan, and Nowé 2013). A more general problem is to optimize the expected value of a potentially nonlinear function of total reward, which may be vector-valued for a multi-objective optimization context. We refer to such a function as a welfare function following (Barman et al. 2023; Fan et al. 2023; Siddique, Weng, and Zimmer 2020); it is also commonly referred to as a scalarization function as in (Hayes et al. 2023; Agarwal, Aggarwal, and Lan 2022).
Most closely related to our work, (Siddique, Weng, and Zimmer 2020) and (Fan et al. 2023) study algorithms with desirable convergenece properties and empirical performance but no approximation guarantees. (Agarwal, Aggarwal, and Lan 2022) describe another model-based algorithm that can compute approximately optimal policies for a general class of monotone and Lipshitz welfare functions, but rather than maximizing the expected welfare, they maximize the welfare of expected rewards (note the two are not equal for nonlinear welfare functions). Other works have formulated fairness in different ways or settings. For example, (Jabbari et al. 2017) defines an analogue of envy freeness and (Deng et al. 2023) studies a per-time-step fairness guarantee. (Barman et al. 2023) studies welfare maximization in multi-armed bandit problems.
To provide provable approximation and runtime guarantees, we return to the classic E3 algorithm (Kearns and Singh 2002). This result showed that general MDPs could be solved to near-optimality efficiently, meaning in time bounded by a polynomial in the size of the MDP (number of states and actions) and the horizon time. Subsequent results refined the achievable bounds (Brafman and Tennenholtz 2003; Auer, Jaksch, and Ortner 2008; Azar, Osband, and Munos 2017).
@@ 3 Preliminaries
We also have a vector-valued reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$. Each of the $d$ dimensions of the reward vector corresponds to one of the multiple objectives that are to be maximized. At each time step $t$, the agent observes state $s_{t}\in\mathcal{S}$, takes action $a_{t}\in\mathcal{A}(s_{t})$, and receives a reward vector $\bm{R}(s_{t},a_{t})\in\mathbb{R}^{d}$. The environment, in turn, transitions into $s_{t+1}$ with probability $Pr(s_{t+1}|a_{t},s_{t})$.
To make the optimization objective well-posed in MOMDPs with vector-valued rewards, we must specify a scalarization function (Hayes et al. 2023) which we denote as $W:{\mathbb{R}}^{d}\to{\mathbb{R}}$. For fair multi-objective reinforcement learning, we think of each of the $d$ dimensions of the reward vector as corresponding to distinct users. The scalarization function can thus be thought of as a welfare function over the users, and the learning agent is a welfare maximizer. Alternatively, even when $d=1$, a nonlinear function $W$ can be a Von Neumann-Morgenstern utility function (Von Neumann and Morgenstern 1947) that expresses the risk-attitudes of the learning agent, with strictly concave functions expressing risk aversion.
Assumptions
Here we state the preliminary assumptions we make about the reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$ and the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$.
  1. 1.
    $\mathbf{R}$ is nonnegative and normalized so that the maximum single-step reward in any dimension is 1.
  2. 2.
    $\mathbf{R}$ is deterministic. Note that the environment remains stochastic due to the transition probabilities. For a model-based approach in scalar-valued or linear reinforcement learning, it is sufficient to simply learn the mean of random reward distributions. This is not true when optimizing the expected value of a nonlinear function. We leave full reward distribution learning to future work.
  3. 3.
    $W$ is monotone non-decreasing. That is, for vectors $\mathbf{y}\geq\mathbf{x}$ with $y_{j}\geq x_{j}$ for $1\leq j\leq d$ , $W(\mathbf{y})\geq W(\mathbf{x})$.
  4. 4.
    $W$ is smooth, specifically, uniformly continuous. For all $\epsilon>0$, $\mathbf{x},\mathbf{y}\geq 0$, there exists $\delta_{\epsilon}$ such that
@@
    Note that uniform continuity is a weaker requirement than Lipshitz continuity.
@@ 4 Modeling Optimality
In this section, we expand the classic model of reinforce- ment learning to optimize the expected value of a nonlinear function of (possibly) multiple dimensions of reward. We begin with the notion of a trajectory of state-action pairs.
@@
Let $M$ be an MOMDP, and let $\pi$ be a policy in $M$. A length $T$ trajectory in $M$ is a tuple $\tau$ of $T$ state-action pairs, that is,
@@
For length $T$ trajectory $\tau$ and discount factor $10\leq\gamma<$ 1, the total discounted reward along $\tau$ is the the vector
$\bm{R}(\tau,\gamma)=\sum_{t=1}^{T}\gamma^{t-1}\bm{R}(s_{t},a_{t}).$
For ease of exposition we will frequently leave $\gamma$ implicit from context and simply write $\bm{R}(\tau)$.
@@
$Pr^{\pi}[\tau]$ $=\pi(a_{1}\mid\tau_{0:0},s_{1})\times$
$\qquad\prod_{t=2}^{T}\pi\left(a_{t}~{}\bigg{|}~{}\tau_{1:t-1},s_{t}\right)Pr(s_{t}|s_{t-1},a_{t-1}).$
Problem Formulation
@@
$\mathbb{E}_{\tau\sim\pi}\bigl{[}W\bigl{(}\bm{R}(\tau)\bigr{)}\bigr{]}=\sum_{\tau}Pr^{\pi}[\tau]W(\mathbf{R}(\tau))$
@@
Note that this objective is not equal to $W\left(\mathbb{E}_{\tau\sim\pi}\bigl{[}\bm{R}(\tau)\bigr{]}\right)$, which others have studied (Agarwal, Aggarwal, and Lan 2022; Siddique, Weng, and Zimmer 2020), for a nonlinear $W$. The former (our objective) is also known as expected scalarized return (ESR) whereas the latter is also known as scalarized expected return (SER) (Hayes et al. 2023). While SER makes sense for a repeated decision-making problem, it does not optimize for expected welfare for any particular trajectory, in contrast to ESR. To the best of our knowledge, our work constitutes the first algorithm with worst-case approximation and polynomial runtime guarantees for ESR.
Form of Optimal Policy and Value Functions
@@
We can similarly write the extended value function of a policy $\pi$ in MORL context. We write $\tau$ as the history or trajectory prior to some current state $s$ and $\tau^{\prime}$ as the future, the remaining $t$ steps determined by the policy $\pi$ and the environmental transitions.
@@
The value of a policy $\pi$ beginning at state $s$ after history $\tau$ and executing $t$ more actions is
@@
Note that because $W$ is nonlinear, the value function $V^{\pi}$ cannot be decomposed in the same way as in the traditional Bellman equations. Instead, we develop an extended form of Bellman optimality by showing that the optimal value function can be written as a function of current state $s$, accumulated discounted reward $\mathbf{R}(\tau)$, and number of timesteps remaining in the task $t$.
@@
Proof.
We proceed by induction on $t$. In the base case of $t=0$, we have simply that $\mathcal{V}(s,\mathbf{R}(\tau),0)=W(\mathbf{R}(\tau))$ by the definition of $\mathcal{V}$. But when $t=0$, the agent is at the terminal state, which implies any $\tau^{\prime}$ in Definition 5 will be the empty trajectory $\tau_{0;0}$. So
For the inductive hypothesis, suppose that the equality holds for all states $s$ and histories $\tau$ for up to $t-1\geq 0$ time steps remaining, i.e.
Then with $t$ steps remaining:
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{\tau^{\prime}}Pr^{\pi}(\tau^{\prime})W(\bm{R}(\tau)+\gamma^{T-t}\bm{R}(\tau^{\prime}))$
where the sum is taken over all length-$t$ trajectories $\tau^{\prime}$ beginning at $s$, and $Pr^{\pi}(\tau^{\prime})$ is the probability that $\tau^{\prime}$ is traversed under policy $\pi$. Note that we can further decompose $\tau^{\prime}$ into
$\tau^{\prime}=(s,a)\oplus\tau^{\prime\prime}$
where $(s,a)$ denotes the state-action pair corresponding to the current timestep, $\oplus$ concatenates, and $\tau^{\prime\prime}$ denotes a trajectory of length $t-1$ beginning at some state $s^{\prime}$ transitioned from state $s$ via taking action $a$. Then we can rewrite $V^{*}$ as
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{a}$ $\pi(a\mid s,\tau,t)~{}\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\bm{R}(\tau)$
$+\gamma^{T-t}\bm{R}(s,a)+\gamma^{T-t+1}\bm{R}(\tau^{\prime\prime}))\bigg{]}$
$\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\bm{R}(\tau)+\gamma^{T-t}\bm{R}(s,a)+\gamma^{T-t+1}\bm{R}(\tau^{\prime\prime}))\bigg{]}$
$V^{*}(s,\tau,t)$ $=\max_{a}\mathbb{E}_{s^{\prime}}\bigg{[}\max_{\pi}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\bm{R}(\tau)$
$\qquad+\gamma^{T-t}\bm{R}(s,a)+\gamma^{T-(t-1)}\bm{R}(\tau^{\prime\prime}))\bigg{]}$
$=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\bm{R}(\tau)+\gamma^{T-t}\bm{R}(s,a),t-1)\right]$
$=\mathcal{V}(s,\bm{R}(\tau),t).$
By Lemma 6, we can parameterize the optimal value function for MORL by the current state $s$, accumulated reward $\mathbf{R_{acc}}\in{\mathbb{R}}^{d}$, and number of timesteps remaining $t$. We will use this formulation of $V^{*}$ in the remainder of the paper.
@@
$V^{*}(s,\mathbf{R}_{acc},t)$ $=\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times$
$\qquad V^{*}(s^{\prime},\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a),t-1).$
Before proceeding we want to provide some intuition as to why the traditional form of Bellman optimality conditions do not hold. The same reasoning helps to explain why stationary policies are not, in general, optimal for our objective.
Horizon Time

### v3

# Multi-objective Reinforcement Learning with Nonlinear Preferences: Provable Approximation for Maximizing Expected Scalarized Return

@@ Abstract
We study multi-objective reinforcement learning with nonlinear preferences over trajectories. That is, we maximize the expected value of a nonlinear function over accumulated rewards (expected scalarized return or ESR) in a multi-objective Markov Decision Process (MOMDP). We derive an extended form of Bellman optimality for nonlinear optimization that explicitly considers time and current accumulated reward. Using this formulation, we describe an approximation algorithm for computing an approximately optimal non-stationary policy in pseudopolynomial time for smooth scalarization functions with a constant number of rewards. We prove the approximation analytically and demonstrate the algorithm experimentally, showing that there can be a substantial gap between the optimal policy computed by our algorithm and alternative baselines.
@@ 1 Introduction
Markov Decision Processes (MDPs) model goal-driven interaction with a stochastic environment, typically aiming to maximize a scalar-valued reward per time-step through a learned policy. Equivalently, this formulation asks the agent to maximize the expected value of a linear function of total reward. This problem can be solved with provable approximation guarantees in polynomial time with respect to the size of the MDP [16, 9, 5, 6, 2].
We extend this framework to optimize a nonlinear function of vector-valued rewards in multi-objective MDPs (MOMDPs), aiming to maximize expected welfare $\mathbb{E}[W(\mathbf{r})]$ where $\mathbf{r}$ is a total reward vector for $d$ objectives. We note that this function is also called the utility or the scalarization function within the multi-objective optimization literature [14]. Unlike deep neural networks used for large state spaces which uses nonlinear function approximators, our nonlinearity lies entirely in the objective function.
Motivation.
@@
Nonlinear welfare functions capture richer preferences for agents, such as fairness or risk attitudes. For example, the Nash Social Welfare function $\left(W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}\right)$ reflects a desire for fairness or balance across objectives and diminishing marginal returns [11]. Even single-objective decision theory uses nonlinear utility functions to model risk preferences, like risk aversion with the Von Neumann-Morgenstern utility function [24].
Consider a minimal example where an autonomous taxi robot, Robbie, serving rides in neighborhoods $A$ and $B$, as diagrammed in Figure 1. Each ride yields a reward in its respective dimension.
$A$$B$$(1,0)$$(0,1)$$(0,0)$
Figure 1: Taxi Optimization Example
Suppose Robbie starts in neighborhood $A$ and has $t=3$ time intervals remaining before recharging. With no discounting, the Pareto Frontier of maximal (that is, undominated) policies can achieve cumulative reward vectors of $(3,0)$, $(1,1)$, or $(0,2)$. If we want Robbie to prefer the second more balanced or “fair” option of serving one ride in each of the two neighborhoods then Robbie must have nonlinear preferences. However, the second option would maximize the Nash Social Welfare of cumulative reward, for example.
Nonlinear preferences complicate policy computation: Bellman optimality fails, and stationary policies may be suboptimal. Intuitively, a learning agent with fairness oriented preferences to balance objectives should behave differently, even in the same state, depending on which dimension of reward is “worse off.” In the Figure 1 example, one policy to achieve the balanced objective $(1,1)$ is to complete a ride in $A$, then travel from $A$ to $B$, and finally to complete a ride in $B$ – note that this is not stationary with respect to the environment states.
Contributions.
In this work, we ask whether it is possible to describe an approximation algorithm (that is, with provable guarantees to approximate the welfare optimal policy) for MOMDPs with nonlinear preferences that has polynomial dependence on the number of states and actions, as is the case for linear preferences or scalar MDPs. To the best of our knowledge, ours is the first work to give provable guarantees for this problem, compared to other work that focuses exclusively on empirical evaluation of various neural network architectures.
We show this is possible for smooth preferences and a constant number of dimensions of reward. To accomplish this, we (i) derive an extended form of Bellman optimality (which may be of independent interest) that characterizes optimal policies for nonlinear preferences over multiple objectives, (ii) describe an algorithm for computing approximately optimal non-stationary policies, (iii) prove the worst-case approximation properties of our algorithm, and (iv) demonstrate empirically that our algorithm can be used in large state spaces to find policies that significantly outperform other baselines.
@@ 2 Related Work
Most algorithms for reinforcement learning focus on a single scalar-valued objective and maximizing total expected reward [21]. Classic results on provable approximation and runtime guarantees for reinforcement learning include the E3 algorithm [16]. This result showed that general MDPs could be solved to near-optimality efficiently, meaning in time bounded by a polynomial in the size of the MDP (number of states and actions) and the horizon time. Subsequent results refined the achievable bounds [9, 5, 6]. We extend these results to the multi-objective case with nonlinear preferences.
Multi-objective reinforcement learning optimizes multiple objectives at once. So-called single-policy methods use a scalarization function to reduce the problem to scalar optimization for a single policy, and we follow this line of research. The simplest form is linear scalarization, applying a weighted sum on the Q vector [18, 17, 1, 25, 4].
A more general problem is to optimize the expected value of a potentially nonlinear function of total reward, which may be vector-valued for a multi-objective optimization context. We refer to such a function as a welfare function [7, 13, 20]; it is also commonly referred to as a utility or scalarization function [14, 3]. Other recent work has studied nonlinear objectives, but ours is the first to our knowledge to provide an approximation algorithm with provable guarantees based on a characterization of recursive optimality in this setting. Several others study algorithms with desirable convergenece properties and empirical performance based on conditioning function approximators on accumulated reward but with no approximation guarantees, all of which we see as complementary efforts to our work [20, 13, 10, 19].
[3] describe another model-based algorithm that can compute approximately optimal policies for a general class of monotone and Lipshitz welfare functions, but rather than maximizing the expected welfare, they maximize the welfare of expected rewards (note the two are not equal for nonlinear welfare functions). Other works have formulated fairness in different ways or settings. For example, [15] defines an analogue of envy freeness and [12] studies a per-time-step fairness guarantee. [7] studies welfare maximization in multi-armed bandit problems.
@@ 3 Preliminaries
We also have a vector-valued reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$. Each of the $d$ dimensions of the reward vector corresponds to one of the multiple objectives that are to be maximized. At each time step $t$, the agent observes state $s_{t}\in\mathcal{S}$, takes action $a_{t}\in\mathcal{A}(s_{t})$, and receives a reward vector $\mathbf{R}(s_{t},a_{t})\in\mathbb{R}^{d}$. The environment, in turn, transitions into $s_{t+1}$ with probability $Pr(s_{t+1}|a_{t},s_{t})$.
To make the optimization objective well-posed in MOMDPs with vector-valued rewards, we must specify a scalarization function [14] which we denote as $W:\mathbb{R}^{d}\to\mathbb{R}$. For fair multi-objective reinforcement learning, we think of each of the $d$ dimensions of the reward vector as corresponding to distinct users. The scalarization function can thus be thought of as a welfare function over the users, and the learning agent is a welfare maximizer. Even when $d=1$, a nonlinear function $W$ can be a Von Neumann-Morgenstern utility function [24] that expresses the risk-attitudes of the learning agent, with strictly concave functions expressing risk aversion.
Assumptions.
Here we clarify some preliminary assumptions we make about the reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$ and the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$. The restriction to $[0,1]$ is not too restrictive and can be seen simply as a normalization; the substantial assumption is that the rewards are simply finite and bounded. Note from the writing we assume the reward function is deterministic: While it suffices for linear optimization to simply learn the mean of random reward distributions, this does not hold when optimizing the expected value of a nonlinear function. Nevertheless, the environment itself may still be stochastic, as the state transitions may still be random.
We also assume that $W$ is smooth. For convenience of analysis, we will assume $W$ is uniformly continuous on the L1-norm (other parameterizations such as the stronger Lipshitz continuity or using the L2-norm are also possible). For all $\epsilon>0$, there exists $\delta_{\epsilon}$ such that for all $\mathbf{x},\mathbf{y}\geq 0$,
@@
This smoothness assumption is necessary for worst-case analysis, but not to run the algorithm.
@@ 4 Modeling Optimality
In this section, we expand the classic model of reinforcement learning to optimize the expected value of a nonlinear function of (possibly) multiple dimensions of reward. We begin with the notion of a trajectory of state-action pairs.
@@
Let $M$ be an MOMDP. A length $T$ trajectory in $M$ is a tuple $\tau$ of $T$ state-action pairs
@@
For length $T$ trajectory $\tau$ and discount factor $10\leq\gamma<$ 1, the total discounted reward along $\tau$ is the vector
$\mathbf{R}(\tau,\gamma)=\sum_{t=1}^{T}\gamma^{t-1}\mathbf{R}(s_{t},a_{t}).$
For ease of exposition we will frequently leave $\gamma$ implicit from context and simply write $\mathbf{R}(\tau)$.
@@
$Pr^{\pi}[\tau]=\pi(a_{1}\mid\tau_{0:0},s_{1})\times\prod_{t=2}^{T}\pi\left(a_{t}~{}\bigg{|}~{}\tau_{1:t-1},s_{t}\right)Pr(s_{t}|s_{t-1},a_{t-1}).$
Problem Formulation.
@@
$\mathbb{E}_{\tau\sim\pi}\bigl{[}W\bigl{(}\mathbf{R}(\tau)\bigr{)}\bigr{]}=\sum_{\tau}Pr^{\pi}[\tau]W(\mathbf{R}(\tau))$
@@
Note that this objective is not equal to $W\left(\mathbb{E}_{\tau\sim\pi}\bigl{[}\mathbf{R}(\tau)\bigr{]}\right)$, which others have studied [3, 20], for a nonlinear $W$. The former (our objective) is also known as expected scalarized return (ESR) whereas the latter is also known as scalarized expected return (SER) [14]. While SER makes sense for a repeated decision-making problem, it does not optimize for expected welfare for any particular trajectory. For concave $W$, ESR $\leq$ SER by Jensen’s inequality. However, an algorithm for approximating SER does not provide any guarantee for approximating ESR. For example, a policy can be optimal on SER but achieve 0 ESR if it achieves high reward on one or the other of two objectives but never both in the same episode.
Form of Optimal Policy and Value Functions.
@@
We can similarly write the extended value function of a policy $\pi$. We write $\tau$ as the history or trajectory prior to some current state $s$ and $\tau^{\prime}$ as the future, the remaining $t$ steps determined by the policy $\pi$ and the environmental transitions.
@@
The value of a policy $\pi$ beginning at state $s$ after history $\tau$ and with $t$ more actions is
@@
Note that because $W$ is nonlinear, the value function $V^{\pi}$ cannot be decomposed in the same way as in the traditional Bellman equations. Before proceeding we want to provide some intuition for this point. The same reasoning helps to explain why stationary policies are not generally optimal.
We develop an extended form of Bellman optimality capturing this intuition by showing that the optimal value function can be written as a function of current state $s$, accumulated discounted reward $\mathbf{R}(\tau)$, and number of timesteps remaining in the task $t$. The proof is deferred to the appendix.
@@
By Lemma 1, we can parameterize the optimal value function by the current state $s$, accumulated reward $\mathbf{R_{acc}}\in\mathbb{R}^{d}$, and number of timesteps remaining $t$. We will use this formulation of $V^{*}$ in the remainder of the paper.
@@
$V^{*}(s,\mathbf{R}_{acc},t)=\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V^{*}(s^{\prime},\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a),t-1).$
Horizon Time.